# On the Complexity of Quantum Field Theory

### v1

@@ 1 Introduction
Quantum field theory (QFT) remains one of the cornerstones of theoretical physics, describing an enormous variety of physical settings with remarkable precision. Though the depth of our understanding of QFTs is ever-increasing, a complete classifications remains elusive. Ultimately, one would like to have a notion of a space of QFTs, in which QFTs and their observables are systematically classified, organized, and relations among them can be understood [1]. An essential part of such a construction is to understand what data fundamentally defines a QFT, and how much information is needed to uniquely specify one. Information-theoretic methods have been widely investigated in the context of QFTs, but currently a precise notion of information content of the data which defines a QFT does not exist.
@@
In both cases, one is naturally led to the fundamental mathematical problem of quantifying the complexity of functions. This is a challenging problem, and evidently impossible for general functions unless some strong finiteness constraints are imposed. For instance, a generic analytic function already requires infinitely many coefficients to be specified in a power series expansion, and thus appears to contain an infinite amount of information. The field of mathematics which studies spaces and functions which contain only a finite amount of information is called tame geometry. Objects in this theory are constrained by a tameness axiom called o-minimality, which essentially demands that they can be constructed with finitely many logical operations [6]. Remarkably, it was found that functions arising in QFTs satisfy this tameness axiom [7, 8, 9, 10], which suggests that they can indeed be described by a finite amount of information [5, 11].
@@
    It is able to describe the complicated functions that appear in QFTs.
@@ 2.1 Complexity of actions – scalar QFTs
There is, however, a fundamental problem with this approach. From the perspective of computational complexity, $N$ and $D$ play a radically different role. This difference is captured by Bézout’s bound, which states that the number of common zeros of $N$ polynomials of degree $D$ is bounded above by $D^{N}$. In particular, it implies that the computational complexity of solving polynomial equations depends polynomially on the degree $D$, whereas it depends exponentially on the number of variables $N$. With this in mind, the complexity of this QFT should reflect the difference in the dependence on $N$ and $D$.
One way to proceed is as follows: instead of combining $N$ and $D$ into a single number, we disentangle the two and measure the complexity by two numbers. At first, this may seem undesirable, since one can then no longer straightforwardly compare the complexities of two different QFTs. In other words, with this proposal there will be no ‘complexity ordering’ on the space of QFTs. However, as we will argue in the remainder of this paper, quantifying complexity by two numbers is actually surprisingly meaningful. Physically, it reflects the fact that a QFT depends on the number of degrees of freedom and the interactions among them in a rather different way. Mathematically, it has deep connections to quantifying complexity in geometry [12, 13], which we explain in more detail in appendix A. For the class of QFTs described by the Lagrangian of equation (2), we could measure the complexity by the pair of integers $(N,D)$. Let us note that, athough this is a good starting point, these integers are not refined enough to capture certain details of the QFT, such as the geometry of the field space and the strength of the coupling constants. We discuss these aspects more carefully later.
@@
The previous discussion now provides a way to describe a large class of QFTs with a finite amount of information, namely those scalar theories for which the potential $V(\phi_{1},\ldots,\phi_{N})$ can be expressed as a Pfaffian function. In particular, it suggests that we measure QFT complexity in a general scalar theory by assigning a format and degree $(\mathcal{F},\mathcal{D})$ to the action. The advantages of this prescription are that it can accurately reflect the number of degrees of freedom and the complexity of interactions, and can describe theories whose actions are formulated with complicated transcendental functions. [^3] When focusing on normalizable theories, it is often sufficient to consider operators of finite dimension in the Lagrangian, but in effective theories much more complicated functions may arise. For example, in the effective actions coming from string theory compactifications one frequently encounters complicated functions coming from period integrals. The number of degrees of freedom is then counted by $\mathcal{F}$, defined as the sum of the number of scalars $N$ and the number of auxiliary functions $r$ appearing in the Lagrangian. The degree $\mathcal{D}$ encodes the complexity of the interactions and relations among the couplings. To gain some intuition for this prescription, let us discuss a few examples.
@@ 2.2 Complexity of actions – general QFTs
For theories with only scalar fields, we have found that the format and degree of the Lagrangian may provide a good measure of QFT complexity. We now extend this discussion to more general theories, including gauge fields, fermions, and non-trivial target spaces. This generalization will lead to a number of conceptual challenges for complexity, and along the way we will discuss how some of these challenges may be resolved.
@@
Counting the number of real field variables, one finds that the complexity of this theory is $(\mathcal{F},\mathcal{D})=\big{(}\tfrac{1}{2}(N^{2}-1)d(d-1),4\big{)}$. Even though the complexity of the theory grows with $N$, it is known that simpler dual descriptions may emerge in large $N$ limits. This phenomenon, where the complexity grows with $N$ but reduces in the strict $N\to\infty$ limit, is the first hint towards an emergence of simplicity captured by the complexity of a QFT, which we will encounter several times in this work.
@@ 2.3 Vacuum structure
Having discussed how to use format and degree to measure the complexity of a QFT action, let us highlight an application. One of the essential properties of format and degree is that they encode bounds on geometric and topological information. For instance, given a function $f$ with complexity $(\mathcal{F},\mathcal{D})$, we can bound the number of connected components of the solution set to the equation $f=0$, analogous to the Bézout bound. The precise form of this bound depends on the class of tame functions under consideration; for the Pfaffian functions we have the bound
@@
When applying this principle to QFTs, we can use this to analyze the vacuum structure of the theory. If the potential $V(\Phi)$ has complexity $(\mathcal{F},\mathcal{D})$, we can then estimate the number of vacua, i.e. the number of connected components of the set
@@
While there is a polynomial growth in $\mathcal{D}$, these estimates grow rapidly with $\mathcal{F}$, and for simple potentials it is much more effective to count the number of vacua by hand. However, for QFTs with extremely complicated potentials, such as the effective theories arising from string compactifications, these countings are numerically very challenging. In such cases, the bounds in terms of format and degree provide simple and universal estimates for the vacuum structure of the theory. [^5] This story holds in general for sharp o-minimality, although one has to adjust the bounds depending on the class of functions under consideration. The framework is designed so that the bounds are always explicitly computable in terms of $(\mathcal{F},\mathcal{D})$.
@@ 2.4 General framework for complexity
A class of functions which has the desired local complexity bounds is given by the log-Noetherian functions introduced in [21]. We will not give the complete definition of these functions (see section 2 of [21]), but rather describe an illustrative case that captures many of their core features. Let us consider a domain of the form
@@
    the domain $U$ is bounded and has to be of a specific cellular form, as explained in detail in [21]. [^6] Essentially, the domains $U$ are constructed inductively by fibering disks and annuli over each other. The radii of these fibered disks and annuli are allowed to vary as a log-Noetherian function on the base space.
A log-Noetherian function is then any function $f(z_{1},\ldots,z_{N})=P(\xi_{1},\ldots,\xi_{R})$ depending polynomially on the functions in the chain. These functions are known to satisfy local complexity bounds, and it has therefore been conjectured that they are sharply o-minimal. While it is presently not known how to precisely assign a format and degree to such a function and domain, it is suggestive from [21] that their sum should be given by
@@
The numbers of zeros of $f$ on $\overline{D_{\circ}(\rho)}$ is given by $1+2\lfloor\frac{\lambda\rho}{2\pi}\rfloor$, and this should be reflected by the complexity of the function. Indeed, in the log-Noetherian chain $\lambda$ appears as a coefficient in the polynomials on the right-hand side, and the dependence on $\rho$ appears through the domain complexity $\mathcal{F}(D_{\circ}(\rho))+\mathcal{D}(D_{\circ}(\rho))$. This shows that the terms in the second line of equation (24) are necessary.
@@
The proposal of [21] is to combine the two constructions explained above; namely to consider the Pfaffian closure of the log-Noetherian functions. This class of functions, though requiring a somewhat complicated construction, turns out to include a wealth of functions, while still retaining controllable local and global complexity bounds. In fact, it is conjectured that this class of functions is sharply o-minimal as well. Notably, the class includes period functions. These are crucial in geometric applications, and also appear ubiquitously in QFTs. For instance, they appear generally in QFT amplitudes through the evaluation of Feynman integrals, and feature strongly in the effective actions obtained from string theory compactifications. We therefore believe that establishing the sharp o-minimality of $\mathbb{R}_{\rm LN,PF}$ bears great significance for exploring the complexity of QFTs.
@@
An important point to emphasize is that sharp o-minimality is defined not only for functions, but also for sets constructed from these functions. In fact, the latter is more fundamental, since functions can be viewed as sets when considering their graphs. The sets considered in sharp o-minimality are built from solution sets of equations, and include for example the domains considered above. The notion of complexity for solution sets of equations should then extend to any collection of sets generated by performing logical operations such as taking finitely many unions, intersections, and projections. In this way, it is possible to assign a format and degree to these sets. Crucially, the framework imposes certain natural requirements on the behavior of format and degree under these logical operations. As an example, the format and degree of sets $A$ and $B$ obey inequalities [^7] Actually, as explained in the appendix, format and degree generate a double filtration on the collection of sets; therefore these inequalities need to be interpreted with some care. such as
@@ 3.1 Amplitudes and correlation functions
Let us first consider amplitudes which are obtained perturbatively by means of Feynman diagrams. We will denote such an amplitude by $\mathcal{A}_{\ell,s}(p,\lambda)$, where $\ell$ is the number of loops, $p=(p_{1},\ldots,p_{n})$ and $s=(s_{1},\ldots,s_{n})$ denote the external momenta and spins, and $\lambda$ collectively denotes the couplings of the theory. In perturbation theory, these amplitudes are computed by summing the values of all Feynman diagrams with $\ell$ loops, and the values of the individual diagrams are obtained by performing certain integrals over rational functions known as Feynman integrals. Recently, it was proven that the resulting functions $\mathcal{A}_{\ell,s}(p,\lambda)$ are tame functions, meaning that they in principle have a finite information content [8]. [^8] The precise statement is that the amplitudes are definable in the o-minimal structure $\mathbb{R}_{\rm an,exp}$, and is based on the definability of the period map [20]. This result opens the door to the functional complexity methods used in the previous subsection. In fact, the amplitudes can be shown to belong to the class of log-Noetherian functions introduced in section 2.4, which is conjectured to be sharply o-minimal in [21].
The complexity of perturbative amplitudes is expected to grow in at least two ways. Firstly, there is a clear dependence on the number of external particles, since the format counts the number of variables. Secondly, at higher orders in perturbation theory one must sum over a larger set of more complicated diagrams, leading to more complicated Feynman integrals and therefore functions of higher complexity.
@@
In terms of functional complexity, the function $f_{j}$ has format and degree $(\mathcal{F},\mathcal{D})=(1,j)$, and in the limit $j\to\infty$, the complexity of $f_{j}$ diverges. However, the limiting function is $f(x)=e^{x}$, which has finite format and degree $(\mathcal{F},\mathcal{D})=(2,2)$ because it admits a simple Pfaffian chain representation as $\partial f/\partial x=f$. Though the information content of the sequence of functions diverges, the limiting function has infinitely many algebraic relations among its coefficients, and therefore a finite complexity. We call this phenomenon the emergence of simplicity.
@@
These can be obtained non-perturbatively through a combination of recursion relations and differential equations [29], and in [5] it was shown that $I_{N}$ has complexity $(\mathcal{F},\mathcal{D})=(4,8+\lceil N/4\rceil)$, growing stepwise linearly with $N$. Following the discussion above, one would then assign to this QFT the complexity class $(\mathcal{F},\mathcal{D})\sim\big{(}O(1),O(N)\big{)}$. Note that the growth of the format and degree have to be quantified separately. In this example, only $\mathcal{D}$ grows, and since computational complexity grows polynomially in $\mathcal{D}$, this indicates that there are algebraic relations present among the observables which reduce their information content.
@@ 3.2 Action to observables
A natural expectation is that theories with an action of higher complexity will have more complicated Feynman rules, and therefore a more complicated perturbation theory. For example, consider scalar QFT with an algebraic scalar potential of complexity $(\mathcal{F},\mathcal{D})$. Every monomial in the potential yields a Feynman rule, and hence there are at most $(\mathcal{F}+\mathcal{D})!/(\mathcal{F}!\mathcal{D}!)$ vertices in the theory. For potentials with only a few interactions, this estimate is sharpened by using the fewnomial representation.

### v2

@@ 1 Introduction
Quantum field theory (QFT) remains one of the cornerstones of theoretical physics, describing an enormous variety of physical settings with remarkable precision. Though the depth of our understanding of QFTs is ever-increasing, a complete classification remains elusive. Ultimately, one would like to have a notion of a space of QFTs, in which QFTs and their observables are systematically classified, organized, and relations among them can be understood [1]. An essential part of such a construction is to understand what data fundamentally defines a QFT, and how much information is needed to uniquely specify one. Information-theoretic methods have been widely investigated in the context of QFTs, but currently a precise notion of information content of the data which defines a QFT does not exist.
@@
In both cases, one is naturally led to the fundamental mathematical problem of quantifying the complexity of functions. This is a challenging problem, and evidently impossible for general functions unless some strong finiteness constraints are imposed. For instance, a generic analytic function already requires infinitely many coefficients to be specified in a power series expansion, and thus appears to contain an infinite amount of information. The field of mathematics which studies spaces and functions which contain only a finite amount of information is called tame geometry. Objects in this theory are constrained by a tameness axiom called o-minimality, which essentially imposes that they have finitely many geometric features [6]. Remarkably, it was found that many functions arising in QFTs satisfy this tameness axiom [7, 8, 9, 10], which suggests that they can indeed be described by a finite amount of information [5, 11].
@@
    It is able to describe many complicated functions that appear in QFTs;
@@ 2.1 Complexity of actions – scalar QFTs
There is, however, a fundamental problem with this approach. From the perspective of computational complexity, $N$ and $D$ play a radically different role. This difference is captured by Bézout’s bound, which states that the number of common zeros of $N$ polynomials of degree $D$ is bounded above by $D^{N}$. In particular, it implies that the computational complexity of solving polynomial equations satisfies a lower bound depending polynomially on the degree $D$, while depending exponentially on the number of variables $N$. With this in mind, the complexity of this QFT should reflect the difference in the dependence on $N$ and $D$.
One way to proceed is as follows: instead of combining $N$ and $D$ into a single number, we disentangle the two and measure the complexity by two numbers. At first, this may seem undesirable, since one can then no longer straightforwardly compare the complexities of two different QFTs. In other words, with this proposal there will be no ‘complexity ordering’ on the space of QFTs. However, as we will argue in the remainder of this paper, quantifying complexity by two numbers is actually surprisingly meaningful. Physically, it reflects the fact that a QFT depends on the number of degrees of freedom and the interactions among them in a rather different way. Mathematically, it has deep connections to quantifying complexity in geometry [12, 13], which we explain in more detail in appendix A. For the class of QFTs described by the Lagrangian of equation (2), we could measure the complexity by the pair of integers $(N,D)$. Let us note that, although this is a good starting point, these integers are not refined enough to capture certain details of the QFT, such as the geometry of the field space and the strength of the coupling constants. We discuss these aspects more carefully later.
@@
The previous discussion now provides a way to describe a large class of QFTs with a finite amount of information, namely those scalar theories for which the potential $V(\phi_{1},\ldots,\phi_{N})$ can be expressed as a Pfaffian function. In particular, it suggests that we measure QFT complexity in a general scalar theory by assigning a format and degree $(\mathcal{F},\mathcal{D})$ to the action. The advantages of this prescription are that it can accurately reflect the number of degrees of freedom and the complexity of interactions, and can describe theories whose actions are formulated with complicated transcendental functions. [^3] When focusing on renormalizable theories, it is often sufficient to consider operators of finite dimension in the Lagrangian, but in effective theories much more complicated functions may arise. For example, in the effective actions coming from string theory compactifications one frequently encounters complicated functions coming from period integrals. The number of degrees of freedom is then counted by $\mathcal{F}$, defined as the sum of the number of scalars $N$ and the number of auxiliary functions $r$ appearing in the Lagrangian. The degree $\mathcal{D}$ encodes the complexity of the interactions and relations among the couplings. To gain some intuition for this prescription, let us discuss a few examples.
@@ 2.2 Complexity of actions – general QFTs
For theories with only scalar fields, we have found that the format and degree of the Lagrangian, defined by means of a Pfaffian chain, may provide a good measure of QFT complexity. We now extend this discussion to more general theories, including gauge fields, fermions, and non-trivial target spaces. This generalization will lead to a number of conceptual challenges for complexity, and along the way we will discuss how some of these challenges may be resolved.
@@
Counting the number of real field variables, one finds that the complexity of this theory is $(\mathcal{F},\mathcal{D})=\big{(}\tfrac{1}{2}(N^{2}-1)d(d-1),2\big{)}$. Even though the complexity of the theory grows with $N$, it is known that simpler dual descriptions may emerge in large $N$ limits. This phenomenon, where the complexity grows with $N$ but reduces in the strict $N\to\infty$ limit, is the first hint towards an emergence of simplicity captured by the complexity of a QFT, which we will encounter several times in this work.
@@ 2.3 Vacuum structure
Having discussed how to use format and degree to measure the complexity of a QFT action, let us highlight an application. One of the essential properties of format and degree is that they encode bounds on geometric and topological information. For instance, given a Pfaffian function $f$ with complexity $(\mathcal{F},\mathcal{D})$, the number of connected components of the solution set to the equation $f=0$ is bounded by [16, 14]
@@
Generalized definitions of format and degree, to be discussed below and in appendix A, satisfy similar bounds growing polynomially in $\mathcal{D}$ [12, 13]. When applying this principle to QFTs, we can use this to analyze the vacuum structure of the theory. If the potential $V(\Phi)$ has complexity $(\mathcal{F},\mathcal{D})$, we can then estimate the number of vacua, i.e. the number of connected components of the set
@@
While there is a polynomial growth in $\mathcal{D}$, these estimates typically grow rapidly with $\mathcal{F}$, and for simple potentials it is much more effective to count the number of vacua by hand. However, for QFTs with extremely complicated potentials, such as the effective theories arising from string compactifications, these countings are numerically very challenging. In such cases, the bounds in terms of format and degree provide simple and universal estimates for the vacuum structure of the theory. [^5] This story holds in general for sharp o-minimality, although one has to adjust the bounds depending on the class of functions under consideration. The framework is designed so that the bounds are always explicitly computable in terms of $(\mathcal{F},\mathcal{D})$.
@@ 2.4 General framework for complexity
A class of functions which satisfies strong local complexity bounds is given by the log-Noetherian functions introduced in [21]. We will not give the complete definition of these functions (see section 2 of [21]), but rather describe an illustrative case that captures many of their core features. Let us consider a domain of the form
@@
    the domain $U$ is bounded and has to be of a specific cellular form, as explained in detail in [21]. [^6] Essentially, the domains $U$ are constructed inductively by fibering disks and annuli over each other. The radii of these fibered disks and annuli are allowed to vary as a log-Noetherian function on the base space. The domain $U$ given in equation (22) is a simple example of such a domain, in which the radii are constant.
A log-Noetherian function is then any function $f(z_{1},\ldots,z_{N})=P(\xi_{1},\ldots,\xi_{R})$ depending polynomially on the functions in the chain. These functions are known to satisfy local complexity bounds, and it has therefore been conjectured that they are sharply o-minimal. While it is presently not known how to precisely assign a format and degree to such a function and domain, it is suggestive [^7] To be precise, it is shown in [21] that these functions are effectively o-minimal, meaning that they satisfy effectively computable bounds characterized by a single number. from [21] that their sum should be
@@
The numbers of zeros of $f$ on $\overline{D_{\circ}(\rho)}$ is given by $2\lfloor\frac{\lambda\rho}{2\pi}\rfloor$, and this should be reflected by the complexity of the function. Indeed, in the log-Noetherian chain $\lambda$ appears as a coefficient in the polynomials on the right-hand side, and the dependence on $\rho$ appears through the domain complexity $\mathcal{F}(D_{\circ}(\rho))+\mathcal{D}(D_{\circ}(\rho))$. This shows that the terms in the second line of equation (24) are necessary.
@@
The proposal of [21] is to combine the two constructions explained above; namely to consider the Pfaffian closure of the log-Noetherian functions. This class of functions, though requiring a somewhat complicated construction, turns out to include a wealth of functions, while still retaining controllable local and global complexity bounds. In fact, it is conjectured that this class of functions is sharply o-minimal as well. Notably, the class includes all algebraic period functions. These are crucial in geometric applications, and also appear ubiquitously in QFTs. For instance, they appear generally in QFT amplitudes through the evaluation of Feynman integrals, and feature strongly in the effective actions obtained from string theory compactifications. We therefore believe that establishing the sharp o-minimality of $\mathbb{R}_{\rm LN,PF}$ bears great significance for exploring the complexity of QFTs.
@@
An important point to emphasize is that sharp o-minimality is defined not only for functions, but also for sets constructed from these functions. In fact, the latter is more fundamental, since functions can be viewed as sets when considering their graphs. The sets considered in sharp o-minimality are built from solution sets of equations, and include for example the domains considered above. The notion of complexity for solution sets of equations should then extend to any collection of sets generated by performing logical operations, i.e. taking finitely many unions, intersections, complements, products, and projections. In this way, it is possible to assign a format and degree to these sets. Crucially, the framework imposes certain natural requirements on the behavior of format and degree under these logical operations. As an example, the format and degree of sets $A$ and $B$ obey inequalities [^8] Actually, as explained in the appendix, format and degree generate a double filtration on the collection of sets; therefore these inequalities need to be interpreted with some care. such as
@@ 3.1 Amplitudes and correlation functions
Let us first consider amplitudes which are obtained perturbatively by means of Feynman diagrams. We will denote such an amplitude by $\mathcal{A}_{\ell,s}(p,\lambda)$, where $\ell$ is the number of loops, $p=(p_{1},\ldots,p_{n})$ and $s=(s_{1},\ldots,s_{n})$ denote the external momenta and spins, and $\lambda$ collectively denotes the couplings of the theory. In perturbation theory, these amplitudes are computed by summing the values of all Feynman diagrams with $\ell$ loops, and the values of the individual diagrams are obtained by performing certain integrals over rational functions known as Feynman integrals. Recently, it was proven that the resulting functions $\mathcal{A}_{\ell,s}(p,\lambda)$ are tame functions, meaning that they in principle have a finite information content [8]. [^9] The precise statement is that the amplitudes are definable in the o-minimal structure $\mathbb{R}_{\rm an,exp}$, and is based on the definability of the period map [20]. This result opens the door to the functional complexity methods used in the previous subsection. In fact, from [8] it follows that the amplitudes belong to the Pfaffian closure of log-Noetherian functions introduced in section 2.4, which is conjectured to be sharply o-minimal in [21].
The complexity of perturbative amplitudes is expected to grow in at least two ways. Firstly, there is a clear dependence on the number of external particles, since the format counts the number of variables. Secondly, at higher orders in perturbation theory one must sum over a larger set of more complicated diagrams, leading to more complicated Feynman integrals and therefore functions with higher $\mathcal{F}$ and $\mathcal{D}$.
@@
In terms of functional complexity, the function $f_{j}$ has format and degree $(\mathcal{F},\mathcal{D})=(1,j)$ when it is viewed as a arbitrary polynomial, and in the limit $j\to\infty$, this complexity diverges. However, these polynomials are not arbitrary, and the limiting function is $f(x)=e^{x}$, which has finite format and degree $(\mathcal{F},\mathcal{D})=(2,2)$ because it admits a simple Pfaffian chain representation as $\partial f/\partial x=f$. The infinitely many algebraic relations among the coefficients ensure that the information content of this sequence of function does not diverge, such that the limiting function has finite complexity. We call this phenomenon the emergence of simplicity.
@@
These can be obtained non-perturbatively through a combination of recursion relations and differential equations [29], and in [5] it was shown that $I_{N}$ has complexity $(\mathcal{F},\mathcal{D})=(4,8+\lceil N/4\rceil)$, growing stepwise linearly with $N$. Following the discussion above, one would then assign to this QFT the complexity class $(\mathcal{F},\mathcal{D})\sim\big{(}O(1),O(N)\big{)}$. Note that the growth of the format and degree have to be quantified separately. In this example, only $\mathcal{D}$ grows, and since computational complexity grows polynomially in $\mathcal{D}$, this indicates that there are algebraic relations present among the observables which reduce their information content. This example strengthens the idea that it is meaningful to quantify complexity by two numbers, as already noted in the mathematical literature.
@@ 3.2 Action to observables
A natural expectation is that theories with an action of higher complexity will have more complicated Feynman rules, and therefore a more complicated perturbation theory. For example, consider a scalar QFT with an algebraic scalar potential of complexity $(\mathcal{F},\mathcal{D})$. Every monomial in the potential yields a Feynman rule, and hence there are at most $(\mathcal{F}+\mathcal{D})!/(\mathcal{F}!\mathcal{D}!)$ vertices in the theory. For potentials with only a few interactions, this estimate is sharpened by using the fewnomial representation.